OpenStudy (anonymous):

Choose an equation in slope-intercept form for the line that passes through (-8,1) and is perpendicular to the y=2x-17. Choose one answer. a. y= -2x+1 b. y= -(1/2)x - 5 c. y= -(1/2)x - 3 d. y= -2x - 4

OpenStudy (anonymous):

@jazy

OpenStudy (anonymous):

Slope of Perpendicular lines = opposite reciprocal ex: Slope of original line = a/b Slope of perpendicular line = -b/a So what would the slope be in the new line?

OpenStudy (anonymous):

hmmm to be honest i dnt even know what that means that u put up there

OpenStudy (anonymous):

Okay , you want a line perpendicular to y=2x-17. The slope of y=2x-17 is 2 or 2/1. Remember the Slope-Intercept form; y = mx + b where m = slope. What is the opposite reciprocal of 2/1? ex: Opposite reciprocal of a/b is -b/a

OpenStudy (anonymous):

2/1 would change into -(1/2). So that leaves you with options b and c.

OpenStudy (anonymous):

its -1/2

OpenStudy (anonymous):

oh lol u said tht

OpenStudy (anonymous):

it was c so thanks

OpenStudy (anonymous):

yw(:

OpenStudy (anonymous):

hay one more question if the 0 is on the bottom like 7/0 is it undefined or 0?

OpenStudy (anonymous):

It is undefined. Remember rise/run. In this case the line doesn't "run"... It only goes straight up and down, "rises". Vertical line... Slope = undefined Horizontal line... Slope = 0